In 2021, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2023. Information related to the contract is as follows: 2021 2022 2023 Cost incurred during the year $2,550,000 $4,250,000 $ 1,870,000 Estimated costs to complete as of year-end 5,950,000 1,700,000 0 Billings during the year 2,050,000 4,750,000 3,200,000 Cash collections during the year 1,825,000 4,100,000 4,075,000 Westgate recognizes revenue over time according to percentage of completion.

Required: 1. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years. (Do not round intermediate calculations. Loss amounts should be indicated with a minus sign.)

2021 2022 2023
Revenue

Gross profit (loss)
Answer #1
  • Requirement 1

2021

2022

2023

Revenue

$       3,000,000.00

$       5,000,000.00

$          2,000,000.00

Gross Profits (Loss)

$           450,000.00

$          750,000.00

$              130,000.00

  • Working

Working

2021

2022

2023

A

Contract Price

$     10,000,000.00

$    10,000,000.00

$        10,000,000.00

B

Cost Incurred to Date

$       2,550,000.00

$       6,800,000.00

$          8,670,000.00

C

Estimated cost yet to be incurred to complete the contract

$       5,950,000.00

$       1,700,000.00

$                                -  

D = B+C

Total Cost

$       8,500,000.00

$       8,500,000.00

$          8,670,000.00

E = (B/D) x 100

% of Completion

30.00%

80.00%

100.00%

F = A x E

Revenue to date

$       3,000,000.00

$       8,000,000.00

$        10,000,000.00

G

Revenue of Previous year

$                            -  

$       3,000,000.00

$          8,000,000.00

H = F - G

Net Revenue this year

$       3,000,000.00

$       5,000,000.00

$          2,000,000.00

I (=B)

Cost to date

$       2,550,000.00

$       6,800,000.00

$          8,670,000.00

J

Cost to date of previous year

$                            -  

$       2,550,000.00

$          6,800,000.00

K = I - J

Net Cost for the year

$       2,550,000.00

$       4,250,000.00

$          1,870,000.00

L = H - K

Gross Profits

$           450,000.00

$          750,000.00

$              130,000.00